the monoatomic ideal gas undergoes a process in which the ratio of P to v at any instant is constant and equals 1. What is the molar heat capacity of the gas?

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
he amount of heat needed to increase the temperature of one mole of a substance by one degree is the molar heat capacity. It is expressed in joules per moles per degrees Celsius (or Kelvin),JoulesMoles∘C. For example, the molar heat capacity of lead is 26.65JoulesMolesK, which means that it takes 26.65 Joules of heat to raise 1 mole of lead by°K.
apply first law of thermodynamics
dQ=dU +PdV
dQ=dU+PdV
dQ/T=Cv+pdV/dT
=3/2R+R/2
=4R/2